### v1

@@ I Introduction
Modules Operations Parameters Input Linear Layer ($L_{\text{input}}$) $(m+1)\times d_{\text{model}}$ Positional Encoding (PE) - Addition ($\text{Add}_{\text{PE}}$) - Encoder Layer MHA Linear Layer ($L_{Q}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ Linear Layer ($L_{K}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ Linear Layer ($L_{V}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ SDA MatMul ($\text{MatMul}_{\text{Score}}$) - Softmax - MatMul ($\text{MatMul}_{\text{Attn}}$) - Output Linear Layer($L_{O}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ Addition ($\text{Add}_{\text{MHA}}$) - BatchNorm ($\text{BN}_{\text{MHA}}$) $2n\times d_{\text{model}}$ FFN Linear Layer ($L_{\text{FFN}_{1}}$) $4(d_{\text{model}}+1)\times d_{\text{model}}$ ReLU - Linear Layer ($L_{\text{FFN}_{2}}$) $(4d_{\text{model}}+1)\times d_{\text{model}}$ Addition ($\text{Add}_{\text{FFN}}$) - BatchNorm ($\text{BN}_{\text{FFN}}$) $2n\times d_{\text{model}}$ Output Global Average Pooling (GAP) - Linear Layer ($L_{\text{output}}$) $d_{\text{model}}+1$
@@ II-A Parameters Simplification
$\text{Params}_{\text{Total}}=12d^{2}_{\text{model}}+(11+m+4n)\times d_{\text{model}}+1$ (1)
@@ V-B Experiments Setup
Each model configuration underwent 50 training sessions consisting of 100 epochs, with early stopping implemented to prevent overfitting. We used a batch size of 256 and the Adam optimizer with standard parameters ($\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=10^{-9}$). The learning rate was initialized at 0.001, with a scheduler having a step size of 3 and a decay factor $\gamma$ of 0.5 for dynamic adjustment during training. The Mean Squared Loss function guided the training process. Post-training, we applied an inverse transformation to the model’s outputs and normalized target values to compute the Root Mean Square Error (RMSE) on the test data as the evaluation metric.
@@
Configs. PeMS AirU Params RMSE Params RMSE n $d_{\text{model}}$ FP32 8-bit 6-bit 4-bit FP32 8-bit 6-bit 4-bit 6 8 1057 0.1609 0.1809 0.2227 0.4224 1105 4.055 4.791 4.696 6.616 16 3649 0.1615 0.1685 0.1988 0.3117 3745 3.726 3.858 4.004 6.566 32 13441 0.1652 0.1563 0.1971 0.3300 13633 3.799 3.665 4.095 5.731 64 51457 0.1666 0.1621 0.1855 0.4084 51841 4.236 3.506 3.923 5.699 12 8 1249 0.1692 0.1847 0.2107 0.4226 1297 3.846 5.274 6.229 6.406 16 4033 0.1582 0.1596 0.1922 0.3979 4129 3.878 4.804 4.675 5.853 32 14209 0.1568 0.1617 0.1848 0.4030 14401 3.859 3.723 4.116 5.474 64 52993 0.1586 0.1571 0.1809 0.3279 53377 4.071 3.518 3.763 5.494 18 8 1441 0.1645 0.2008 0.2158 0.5258 1489 3.917 5.419 6.757 6.765 16 4417 0.1593 0.1607 0.1838 0.3297 4513 3.662 4.974 4.614 5.866 32 14977 0.1567 0.1581 0.1918 0.3162 15169 3.603 3.796 4.012 5.795 64 54529 0.1600 0.1583 0.1845 0.3462 54913 4.055 3.518 3.897 5.286 24 8 1633 0.1664 0.2229 0.2436 0.4841 1681 4.129 5.377 7.160 6.795 16 4801 0.1595 0.1673 0.1988 0.3856 4897 3.899 4.929 5.776 5.881 32 15745 0.1574 0.1591 0.1943 0.4158 15937 3.867 4.019 4.380 5.700 64 56065 0.1604 0.1573 0.1874 0.3787 56449 3.988 3.619 3.840 5.456
@@ V-E Timing Analysis
Focusing on the three candidate configurations identified in Section V-D, we analyzed the timing performance of these FPGA-deployable models. As detailed in the third column of Table IV, the clock frequency of the 8-bit quantized model was limited to 100 MHz. In contrast, models with lower bitwidths could operate at frequencies up to 25% higher, supporting the expectation that fewer bits expedite computation due to simplified logic. However, decreasing the bitwidth from 6-bit to 4-bit did not further increase the frequency, likely due to reduced DSPs engagement and subsequent increases in logic delay.

### v4

@@ I Introduction
Modules Operations Parameters Input Linear Layer ($L_{\text{input}}$) $(m+1)\times d_{\text{model}}$ Positional Encoding (PE) - Addition ($\text{Add}_{\text{PE}}$) - Encoder Layer MHA Linear Layer ($L_{Q}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ Linear Layer ($L_{K}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ Linear Layer ($L_{V}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ SDA MatMul ($\text{MatMul}_{\text{Score}}$) - Softmax - MatMul ($\text{MatMul}_{\text{Attn}}$) - Output Linear Layer($L_{O}$) $(d_{\text{model}}+1)\times d_{\text{model}}$ Addition ($\text{Add}_{\text{MHA}}$) - BatchNorm ($\text{BN}_{\text{MHA}}$) $2d_{\text{model}}$ FFN Linear Layer ($L_{\text{FFN}_{1}}$) $4(d_{\text{model}}+1)\times d_{\text{model}}$ ReLU - Linear Layer ($L_{\text{FFN}_{2}}$) $(4d_{\text{model}}+1)\times d_{\text{model}}$ Addition ($\text{Add}_{\text{FFN}}$) - BatchNorm ($\text{BN}_{\text{FFN}}$) $2d_{\text{model}}$ Output Global Average Pooling (GAP) - Linear Layer ($L_{\text{output}}$) $d_{\text{model}}+1$
@@ II-A Parameters Simplification
$\text{Params}_{\text{Total}}=12d^{2}_{\text{model}}+(15+m)\times d_{\text{model}}+1$ (1)
@@ V-B Experiments Setup
Each model configuration underwent 50 training sessions consisting of 100 epochs, with early stopping implemented to prevent overfitting. We used a batch size of 256 and the Adam optimizer with standard parameters ($\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.98$, $\epsilon\!=\!10^{-9}$). The learning rate was initialized at 0.001, with a scheduler having a step size of 3 and a decay factor $\gamma$ of 0.5 for dynamic adjustment during training. These training sessions were conducted on an NVIDIA GeForce RTX 2080 SUPER GPU, utilizing CUDA 11.0 and PyTorch 3.11 within the Ubuntu 20.04.6 LTS. The objective metric for training was the minimization of the Mean Squared Error. The Mean Squared Loss function guided the training process. Post-training, we applied an inverse transformation to the model’s outputs and normalized target values to compute the Root Mean Square Error (RMSE) on the test data as the evaluation metric.
@@
Configs. PeMS AirU Params RMSE Params RMSE n $d_{\text{model}}$ FP32 8-bit 6-bit 4-bit FP32 8-bit 6-bit 4-bit 6 8 897 0.1609 0.1809 0.2227 0.4224 945 4.055 4.791 4.696 6.616 16 3329 0.1615 0.1685 0.1988 0.3117 3425 3.726 3.858 4.004 6.566 32 12801 0.1652 0.1563 0.1971 0.3300 12993 3.799 3.665 4.095 5.731 64 50177 0.1666 0.1621 0.1855 0.4084 50561 4.236 3.506 3.923 5.699 12 8 897 0.1692 0.1847 0.2107 0.4226 945 3.846 5.274 6.229 6.406 16 3329 0.1582 0.1596 0.1922 0.3979 3425 3.878 4.804 4.675 5.853 32 12801 0.1568 0.1617 0.1848 0.4030 12993 3.859 3.723 4.116 5.474 64 50177 0.1586 0.1571 0.1809 0.3279 50561 4.071 3.518 3.763 5.494 18 8 897 0.1645 0.2008 0.2158 0.5258 945 3.917 5.419 6.757 6.765 16 3329 0.1593 0.1607 0.1838 0.3297 3425 3.662 4.974 4.614 5.866 32 12801 0.1567 0.1581 0.1918 0.3162 12993 3.603 3.796 4.012 5.795 64 50177 0.1600 0.1583 0.1845 0.3462 50561 4.055 3.518 3.897 5.286 24 8 897 0.1664 0.2229 0.2436 0.4841 945 4.129 5.377 7.160 6.795 16 3329 0.1595 0.1673 0.1988 0.3856 3425 3.899 4.929 5.776 5.881 32 12801 0.1574 0.1591 0.1943 0.4158 12993 3.867 4.019 4.380 5.700 64 50177 0.1604 0.1573 0.1874 0.3787 50561 3.988 3.619 3.840 5.456
@@ V-E Timing Analysis
Focusing on the three candidate configurations identified in Section V-D, we analyzed the timing performance of these FPGA-deployable models. As detailed in the third column of Table IV, the clock frequency of the 8-bit quantized model was limited to 100 MHz. In contrast, models with lower bitwidths operate at frequencies up to 25% higher, supporting the expectation that fewer bits expedite computation due to simplified logic. However, decreasing the bitwidth from 6-bit to 4-bit did not increase the frequency, likely due to reduced DSPs engagement and subsequent increases in logic delay.